OpenStudy (anonymous):

Solve 2(4 – b) = 2 − b –(b − 6).

mathslover (mathslover):

\[\huge{8-2b=2-b-b+6}\] \[\huge{8-2b=2-2b+6}\] \[\huge{0=0}\] lolz ... the equation is wrong or b HAS NO SOLUTIONS

OpenStudy (anonymous):

That was one of my answers haha! Thank you!